# A Framework to Audit Email Address Privacy and Analyze Email Marketing Practices of Online Services and Apps

### v1

@@ Abstract.
This study explores the widespread perception that personal data, such as email addresses, may be shared or sold without informed user consent, investigating whether these concerns are reflected in actual practices of popular online services and apps. Over the course of a year, we collected and analyzed the source, volume, frequency, and content of emails received by users after signing up for the 150 most popular online services and apps across various sectors. By examining patterns in email communications, we aim to identify consistent strategies used across industries, including potential signs of third-party data sharing. This analysis provides a critical evaluation of how email marketing tactics may intersect with data-sharing practices, with important implications for consumer privacy and regulatory oversight. Our study findings, conducted post-CCPA and GDPR, indicate that while no third-party spam email was detected, internal email marketing practices were pervasive, with companies frequently sending promotional and CRM emails despite opt-out preferences. The framework established in this work is designed to be scalable, allowing for continuous monitoring, and can be extended to include a more diverse set of apps and services for broader analysis, ultimately contributing to improved user perception of data privacy practices.
conference: ; ;
@@ 1. Introduction
Public perceptions of data privacy, particularly concerning personal data such as email addresses, can be significantly shaped by high-profile data breaches, feelings of powerlessness, and general misunderstandings. Isolated incidents of major data breaches contribute to the growing concern among users about the security of their personal information. For instance, the Yahoo data breach in 2013–2014 affected all 3 billion user accounts, exposing email addresses and other personal data (New York Times, 2017a). Similarly, the Equifax breach in 2017 compromised the personal information of approximately 147 million people, including email addresses and Social Security numbers (New York Times, 2017b). These incidents have heightened public awareness of vulnerabilities in corporate data handling practices. Such breaches exacerbate the concept of “digital resignation” where individuals feel powerless to control their personal information, accepting that their email addresses and other data may be shared or exposed without their consent (Draper and Turow, 2019). This resignation is further fueled by past corporate practices that shared data with third parties (Englehardt et al., 2018), before data privacy laws such as GDPR (Voigt and von dem Bussche, 2017) and CCPA (Pardau, 2018) were enacted. Hoofnagle et al. (Hoofnagle et al., 2012) found that users are often unaware of the extent to which their email addresses are collected and shared with third-parties, increasing feelings of vulnerability and mistrust. Furthermore, Acquisti et al. (Acquisti et al., 2015) found that even users who express privacy concerns can themselves engage in behaviors that compromise their email data, often due to a lack of viable alternatives or understanding.
@@
A significant challenge for users is the difficulty of identifying which app or service may have leaked or shared their email address when they start receiving unsolicited spam or phishing emails. This issue arises because most users generally reuse the same email address across multiple online services and apps. Consequently, when unsolicited communications are received, users cannot determine the specific source responsible for the unauthorized sharing or leakage of their email information. This difficulty in identifying the source of leaked or shared email addresses is further exacerbated by pervasive email marketing practices, which were often connected to the sharing of email and other personal data with third parties. Companies often engaged in email marketing strategies that involve partnering with affiliates, utilizing data brokers, or participating in advertising networks to broaden their outreach (Sherman, 2021). These practices frequently entailed sharing or selling users’ email addresses, sometimes without informed consent or clear disclosure. As a result, users could receive unsolicited promotional emails from entities with whom they have no direct relationship, obscuring the boundaries of accountability. The interconnection between email marketing and email sharing practices not only increases the volume of unsolicited communications but also undermines users’ ability to trust that their personal information is handled responsibly.
@@
After the introduction of data privacy laws like GDPR and CCPA, there still remains a significant gap towards systematically auditing online services and apps for their email address sharing and email marketing practices. Email marketing strategies have become increasingly sophisticated, often blurring the lines between legitimate marketing and unauthorized data sharing with third parties. Without robust auditing mechanisms, it’s challenging to assess compliance with these regulations or hold companies accountable for improper handling of email data. Developing such a framework is crucial for enhancing transparency and accountability in email marketing practices, ultimately improving user perceptions by rebuilding trust in how their personal information is managed.
@@
Towards addressing this gap, our study offers a structured and scalable framework to assess how online services and apps handle user data. Our contributions include the following key steps and insights:
@@
  • Audit of Online Services and Apps: We conducted a year-long audit of the top 150 popular online services and apps, opting out of any optional communication during the registration process. No third-party spam emails were detected within this time frame, suggesting that opting out during account creation is very effective in limiting unsolicited emails.
@@
  • Email Marketing Analysis: Despite no unsolicited emails, we collected 4,847 emails received directly from these online services and apps. Our analysis revealed that companies still frequently used customer relationship management (CRM), alert, and promotional emails to engage users, even when opting out of promotional communications. This highlights the persistence of well-established marketing strategies, implemented internally.
@@
  • Cross-Sector Insights: We identified significant differences in email marketing practices across sectors. Consumer industries, such as retail, heavily relied on promotional campaigns, while financial and service sectors prioritized CRM and alert emails. Some of these sector-specific behaviors offer insights into varying risks to data privacy.
@@
These findings contribute to a comprehensive understanding of email marketing and data sharing practices, promoting transparency and accountability in how online services and apps handle user data.
@@ 2. Related Work
Email privacy has become a critical issue in the wake of increased data breaches and unauthorized sharing of personal information. Englehardt et al. (Englehardt et al., 2018) demonstrated the widespread use of email tracking, allowing third parties to gather user behavior data without consent. Their findings emphasize how email addresses, once shared, become tools for covert tracking, often leading to privacy violations. Similarly, Hoofnagle et al. (Hoofnagle et al., 2012) found that most users are unaware of how their email data is shared with third parties for advertising purposes, heightening concerns about privacy infringement. Plice et al.  (Plice et al., 2008) provided an economic analysis of unwanted commercial messages and their dependence on weak privacy controls. These studies highlight the necessity of open auditing mechanisms to protect email privacy.
@@
Existing literature highlights the need for a balance between effective email marketing and safeguarding email privacy. While predictive models and personalized strategies can optimize engagement, they must be deployed ethically to avoid violating user privacy. Prior works emphasize that compliance with data protection laws, along with the adoption of transparent email marketing practices, is critical for maintaining consumer trust. Our study builds on this foundation by analyzing the email marketing and data-sharing practices of popular apps and services. Unlike prior works, which often did in-depth analysis of emails pertaining to a single company or industry type, our study instead focuses on a breath of services and apps across various sectors.
3. Experimental Setup and Data Collection
Our experimental setup is based on the assumption that if email addresses are leaked, shared, or sold by the services or apps we signed up for, those addresses would eventually start receiving spam or unsolicited emails from third parties. This assumption allows us to detect unauthorized sharing of user data through the presence of unsolicited communications, which are typically linked to marketing activities that involve purchasing user data from third-party vendors. It is important to note, however, that this approach has limitations. Specifically, we cannot detect instances where email addresses are leaked, shared or sold if those third parties do not subsequently send spam or other detectable communications.
To comprehensively evaluate the email sharing and marketing practices across different sectors, we selected a diverse set of widely used online services, mobile applications, and essential services, as follows:
@@ 3. Experimental Setup and Data Collection
This system facilitated clear differentiation between service types and ensured that each service and app had a unique identifier throughout the experiment. During registration, no additional subscriptions, email notifications, or promotional materials were selected. This was essential for maintaining consistency in data collection, allowing us to focus on analyzing unsolicited marketing emails from the service and app providers.
@@
Email Content Classification. To classify each email into one of the three classes (promotional, CRM, or alert), we employed Llama3.1, a leading, open source, large language model (LLM) given definitions directly from prior work (Thomas et al., 2022a). We utilized an in-context learning approach for this task, where the context includes the task description, the definitions from the prior work (the context), and the instance to be classified (Loukas et al., 2023). This approach has been shown to achieve human-level performance on sufficiently large LLMs with clear and concise rules (Brown et al., 2020). The full procedure is available in Appendix A. Understanding the distribution of these categories provides insights into how businesses strategically balance promotional campaigns, customer relationship management, and essential service notifications.
@@ 4.1. Descriptive Statistics
Over the 361-day collection period, we received 4,847 emails from 111 of the 150 services and apps, post registration (LABEL:table:all-domains in Appendix B). Of those emails received, 99.96% passed SPF checks and 81.64% passed DKIM checks. All emails that passed DKIM checks also passed SPF checks. Sender Policy Framework (SPF) verifies the sender’s IP address to ensure it is authorized by the domain owner, while DomainKeys Identified Mail (DKIM) adds a cryptographic signature to emails, verifying that the message hasn’t been altered in transit and is indeed from the claimed sender. Both mechanisms are crucial for protecting against spoofing and phishing attacks, which is why their high pass rates are significant in confirming the authenticity of the emails in our dataset. Post inspection of the emails received by their root domains and the email security checks via SPF and DKIM, we confirmed that the emails received were from the original companies’ domains with which we had made accounts and no third-party domain emails were detected in this dataset.
Initial descriptive statistics indicate that there is not a well-defined parametric distribution that determines the volume of emails received. The mean was 43 emails per domain with a median of 5, a skewness of 3.55 showing the heavy right tail, and kurtosis of 12.92. The data’s high skewness and kurtosis suggest a complex, highly skewed, and heavy-tailed nature that is not well captured by parametric distributions. Figure 1 shows that a few large senders make up the dominant amount of email volume received, with the cumulative distribution following the 80-20 Pareto principle, where the total volume of emails received are dominated by a few senders. Just the top 10 root domains make up 63.23% of the total volume.
@@
To classify the content of the emails, we rely on definitions provided by (Thomas et al., 2022a), where emails were classified as either “promotional”, “CRM” (customer relationship management), or “alerts”. These email types convey distinct messages with differing levels of persuasion. It is the appropriate combination of these different types of messages at the right time that drive click-through and eventual conversion, or conversely, abandonment, by the potential customer. To classify the emails that we collected, we used an in-context classification approach with a general purpose LLM (Llama3.1-8b-instruct (AI@Meta, 2024)) with structured output. The prompt and structured output schema is provided in Appendix A. There were 3,346 (69.0%) promotional emails, 1,393 (28.7%) CRM emails, 92 (1.89%) alert emails, and 16 (0.33%) that were unable to be classified as they were unparseable.
@@ 4.2. Senders by Autonomous System
We captured the IP address of the sender of each email, with that we then ran a lookup of the registered autonomous system for each of those IP addresses and were able to tie it back to the company that sent it. Because of this, we are able to identify that most of the emails received come from a third-party intermediary that specializes in business-to-business marketing and CRM services like Salesforce and Mailgun. Like the overall email volume, the email volume by ASN also follows the Pareto principle, where a few large providers are responsible for the sending of the large majority of emails. Specifically, the top twenty percent, 8 ASNs, produced 89.35% of the total email volume. Figure 2 shows the evidence that most of the email marketing flow originates from a few large, third-party providers.
@@
However, understanding which companies use which email marketing service, or whether they use a third-party service at all, is not straightforward. Some companies send emails originating from their own registered ASN such as wish.com (Wish is the “doing-business-as” name for ContextLogic), Alibaba Group, LinkedIn, and Adobe (Adobe acquired Omniture as part of their web analytics platform). Others use specialized third parties, chiefly Salesforce, Mailgun, Sendgrid, and Sparkpost. A third group had emails originate from cloud service providers like AWS, GCP, and Rackspace. For those that originate from cloud service providers, using only the IP addresses we cannot ascertain what entity is managing the email campaigns at the application layer.

### v2

# Why You’ve Got Mail: Evaluating Inbox Privacy Implications of Email Marketing Practices in Online Apps and Services

@@ Abstract.
This study explores the widespread perception that personal data, such as email addresses, may be shared or sold without informed user consent, investigating whether these concerns are reflected in actual practices of popular online services and apps. Over the course of a year, we collected and analyzed the source, volume, frequency, and content of emails received by users after signing up for the 150 most popular online services and apps across various sectors. By examining patterns in email communications, we aim to identify consistent strategies used across industries, including potential signs of third-party data sharing. This analysis provides a critical evaluation of how email marketing tactics may intersect with data-sharing practices, with important implications for consumer privacy and regulatory oversight. Our study findings, conducted post-CCPA and GDPR, indicate that while no unknown third-party spam email was detected, internal and authorized third-party email marketing practices were pervasive, with companies frequently sending promotional and CRM emails despite opt-out preferences. The framework established in this work is designed to be scalable, allowing for continuous monitoring, and can be extended to include a more diverse set of apps and services for broader analysis, ultimately contributing to transparency in email address privacy practices.
Email Privacy, Email Marketing, Unsolicited Email, Spam.
journalyear: 2025conference: Proceedings of the ACM Conference on Data and Application Security and Privacy 2025; June 4–June 6, 2025; Pittsburgh, PA, USAbooktitle: Proceedings of the ACM Conference on Data and Application Security and Privacy 2025 (CODASPY ’25), June 4–June 6, 2025, Pittsburgh, PA, USA
Taxonomies
We distinguish email sources based the following classifications:
  • Internal: Emails sent directly by the service’s own infrastructure. They typically should pass Sender Policy Framework (SPF) and DomainKeys Identified Mail (DKIM) checks.
  • Authorized Third-Party (ATP): Emails sent on behalf of the service by authorized providers, authenticated by the service. They typically should also pass SPF and DKIM checks.
  • Unknown Third-Party (UTP): Emails from unauthorized or unknown entities. If such an email is masqueraded as coming from the service, it typically will fail SPF and DKIM checks.
Unsolicited emails are grouped into these broad categories:
  • Service-Originated Spam (SOS): Unsolicited internal and ATP emails that pass SPF and DKIM checks.
  • Unknown or Unauthenticated Sender Spam (UUSS): Emails from UTP sources and emails that fail SPF and DKIM checks (even if the header indicates service-originated in the From: field).
Emails are further categorized by purpose:
  • CRM: Customer Relationship Management emails aimed at engagement and relationship-building.
  • Promotional: Emails with overt selling intent, such as discounts or offers.
  • Alerts: Informational emails, such as account notifications or updates.
@@ 1. Introduction
Public perceptions of data privacy, particularly concerning personal data such as email addresses, can be significantly shaped by high-profile data breaches, feelings of powerlessness, and general misunderstandings. Isolated incidents of major data breaches contribute to the growing concern among users about the security of their personal information. For instance, the Yahoo data breach in 2013–2014 affected all 3 billion user accounts, exposing email addresses and other personal data (New York Times, 2017a). Similarly, the Equifax breach in 2017 compromised the personal information of approximately 147 million people, including email addresses and Social Security numbers (New York Times, 2017b). These incidents have heightened public awareness of vulnerabilities in corporate data handling practices. Such breaches exacerbate the concept of “digital resignation” where individuals feel powerless to control their personal information, accepting that their email addresses and other data may be shared or exposed without their consent (Draper and Turow, 2019). This resignation is further fueled by past corporate practices that shared data with third-parties (Englehardt et al., 2018), before data privacy laws such as GDPR (Voigt and von dem Bussche, 2017) and CCPA (Pardau, 2018) were enacted. Hoofnagle et al. (Hoofnagle et al., 2012) found that users are often unaware of the extent to which their email addresses are collected and shared with third-parties, increasing feelings of vulnerability and mistrust. Furthermore, Acquisti et al. (Acquisti et al., 2015) found that even users who express privacy concerns can themselves engage in behaviors that compromise their email data, often due to a lack of viable alternatives or understanding.
@@
A significant challenge for users is the difficulty of identifying which app or service may have leaked or shared their email address when they start receiving unsolicited spam or phishing emails. This issue arises because most users generally reuse the same email address across multiple online services and apps. Consequently, when unsolicited communications are received from third-parties, users may not be able to determine the specific source responsible for any unauthorized sharing or leakage of their email information. This difficulty in identifying the source of leaked or shared email addresses is further exacerbated by pervasive email marketing practices, which were often connected to the sharing of email and other personal data with third-parties. Companies often engaged in email marketing strategies that involve partnering with affiliates, utilizing data brokers, or participating in advertising networks to broaden their outreach (Sherman, 2021). These practices frequently entailed sharing or selling users’ email addresses, sometimes without informed consent or clear disclosure. As a result, users could receive unsolicited promotional emails from entities with whom they have no direct relationship, obscuring the boundaries of accountability. The interconnection between email marketing and email sharing practices not only increases the burden of unsolicited communications (Vacek, 2014; Agema, 2015) but also undermines users’ ability to trust that their personal information is handled responsibly.
@@
After the introduction of data privacy laws like GDPR and CCPA, there still remains a significant gap towards systematically auditing online services and apps for their email address sharing and email marketing practices. Email marketing strategies have become increasingly sophisticated, often blurring the lines between legitimate marketing and unauthorized data sharing with third-parties. Without robust auditing mechanisms, it’s challenging to assess compliance with these regulations, hold companies accountable for improper handling of email data, or impose consequences for sending unsolicited emails despite user opt-outs. Developing such a framework is crucial for enhancing transparency and accountability in email marketing practices, ultimately improving user perceptions by rebuilding trust in how their personal information is managed.
@@
Towards addressing this gap, our study offers a structured and scalable framework to assess how online services and apps handle users’ email addresses, from the perspective of an user inbox. Our contributions include the following key steps and insights:
@@
  • Audit of Online Services and Apps: We conducted a year-long audit of email communications received after registering with the top 150 popular online services and apps, opting out of any optional communication during the registration process. No unknown or unauthenticated sender spam (UUSS) emails were detected within this time frame, suggesting that opting out during account creation is very effective in limiting unsolicited emails from unknown and unauthenticated entities.
@@
  • Email Marketing Analysis: Despite no UUSS emails, we collected 4,847 emails received from these online services and apps, directly (internal) or from authorized third-parties (ATP). Our analysis revealed that companies still frequently used customer relationship management (CRM), alert, and promotional emails to engage users, even when opting out of promotional communications. This highlights the persistence of well-established marketing strategies, majority of which may be classified as service-originated spam (SOS).
@@
  • Cross-Sector Insights: We identified significant differences in email marketing practices across sectors. Consumer industries, such as omnichannel retail, heavily relied on promotional campaigns, while financial services prioritized CRM and alert emails. Some of these sector-specific behaviors offer insights into varying risks to inbox privacy.
@@
These findings contribute to a broader understanding of email marketing and data sharing practices, promoting transparency and accountability in how online services and apps handle users’ email addresses.
Figure 1. Framework for email collection, processing, and classification to analyze inbox privacy across services.
@@ 2. Related Work
Email and inbox privacy has become a critical issue in the wake of increased data breaches and unauthorized sharing of personal information. Englehardt et al. (Englehardt et al., 2018) demonstrated the widespread use of email tracking, allowing third parties to gather user behavior data without consent. Their findings emphasize how email addresses, once shared, become tools for covert tracking, often leading to privacy violations. Similarly, Hoofnagle et al. (Hoofnagle et al., 2012) found that most users are unaware of how their email data is shared with third parties for advertising purposes, heightening concerns about privacy infringement. Plice et al.  (Plice et al., 2008) provided an economic analysis of unwanted commercial messages and their dependence on weak privacy controls. These studies highlight the necessity of open auditing mechanisms to protect email privacy.
@@
Existing literature highlights the need for a balance between effective email marketing and safeguarding inbox privacy. While predictive models and personalized strategies can optimize engagement, they must be deployed ethically to avoid violating user privacy. Prior works emphasize that compliance with data protection laws, along with the adoption of transparent email marketing practices, is critical for maintaining consumer trust. Our study builds on this foundation by analyzing the email marketing and data-sharing practices of popular apps and services. Unlike prior works, which often did in-depth analysis of emails pertaining to a single company or industry type, our study instead focuses on a breadth of services and apps across various sectors.
3. Data Collection Framework
Our framework (Figure 1) is based on the assumption that if email addresses are leaked, shared, or sold by the services or apps we signed up for, those addresses would eventually start receiving spam or unsolicited emails from third-parties. This assumption allows us to passively detect unauthorized sharing of user data through the presence of unsolicited communications, which are typically linked to marketing activities that involve purchasing user data from third-party vendors. It is important to note, however, that this approach has its limitations. Specifically, we cannot detect instances where email addresses are leaked, shared or sold to third-parties if they do not subsequently send spam or other detectable communications.
To comprehensively evaluate the email sharing and marketing practices across different sectors, we selected a diverse set of widely used online services and mobile applications, as follows:
@@ 3. Data Collection Framework
This design of our framework facilitates clear differentiation between service types and ensures that each service and app has a unique identifier throughout the experiment. During registration, no additional subscriptions, email notifications, or promotional materials were selected. This was essential for maintaining consistency in data collection, allowing us to focus on analyzing unsolicited marketing emails from the service and app providers.
@@
Email Content Classification. To classify the content of the emails, we rely on definitions provided by Thomas et al. (Thomas et al., 2022a), where emails were classified as either promotional, CRM, or alerts. These email types convey distinct messages with differing levels of persuasion. It is the appropriate combination of these different types of messages at the right time that drive click-through and eventual conversion, or conversely, abandonment, by the potential customer. To classify the emails that we collected, we used an in-context classification approach with a general purpose LLM (Llama3.1-8b-instruct (AI@Meta, 2024)) with structured output. This approach has been shown to achieve human-level performance on sufficiently large LLMs with clear and concise rules (Brown et al., 2020). The full procedure is available in Appendix A. There were 3,346 (69.0%) promotional emails, 1,393 (28.7%) CRM emails, 92 (1.89%) alert emails, and 16 (0.33%) that were unable to be classified as they were unparseable.
The use of LLMs as a rater for content classification tasks such as this is an area of active research (Theelen et al., 2024; Chae and Davidson, 2023), and recently has been successfully applied to spam email classification tasks (Rojas-Galeano, 2024). For this task, however, which relies on the LLM to perform comparably to a human rater, we enlisted 5 independent researchers from our institution to classify 50 randomly selected emails using the same instructions given to the LLM. Since there is no ground-truth, we instead evaluate whether the LLM-human inter-rater reliability (IRR) falls within the range of human-human IRR. The average Cohen’s Kappa among human-human pairs is $\kappa=0.406$ (moderate agreement), and the average LLM-human Kappa is $\kappa=0.479$ (moderate agreement). In other words, the LLM achieves a level of agreement with humans that not only exceeds the human-human average, but also places it well within the variability observed among human raters themselves. This suggests that the LLM is performing comparably to human raters, demonstrating its viability as a reasonable evaluator in this classification task.
Our framework and following analyses address key aspects of data privacy laws, including the right to opt-out of marketing communications and restrictions on unauthorized data sharing practices, and can be extended to evaluate compliance across more diverse regulatory environments.
@@ 4.1. Descriptive Statistics
Over the 361-day collection period, we received 4,847 emails from 109 of the 150 services and apps, post registration (LABEL:table:all-domains in Appendix B). Of those emails received, 99.96% passed SPF checks and 81.64% passed DKIM checks. All emails that passed DKIM checks also passed SPF checks. Sender Policy Framework (SPF) verifies the sender’s IP address to ensure it is authorized by the domain owner, while DomainKeys Identified Mail (DKIM) adds a cryptographic signature to emails, verifying that the message hasn’t been altered in transit and is indeed from the claimed sender. Although ideally, both SPF and DKIM should pass, the successful verification of either check indicates that the email was authorized by the service with high confidence. A detailed manual review of the three emails that failed both SPF and DKIM revealed that these failures were likely due to DNS or server misconfigurations rather than malicious intent.
Post inspection of the emails received by their root domains and the email security checks via SPF and DKIM, we conclude that all emails received were from the original companies’ domains with which we had made accounts and no UUSS emails were detected in this dataset.
However, further analysis reveals that the origin of the emails tell a more complex story of the lineage and custody of user data. Initial descriptive statistics indicate that there is not a well-defined parametric distribution that determines the volume of emails received. The mean was 43 emails per domain with a median of 5, a skewness of 3.55 showing the heavy right tail, and kurtosis of 12.92. The data’s high skewness and kurtosis suggest a complex, highly skewed, and heavy-tailed nature that is not well captured by parametric distributions. Figure 2 shows that a few large senders make up the dominant amount of email volume received, with the cumulative distribution following the 80-20 Pareto principle, where the total volume of emails received are dominated by a few senders. Just the top 10 root domains make up 63.23% of the total volume.
@@ 4.2. Senders by Autonomous System
We captured the IP address of the sender of each email, with that we then ran a lookup of the registered autonomous system for each of those IP addresses and were able to tie it back to the company that sent it. Because of this, we are able to identify that most of the emails received come from authorized third-parties that specializes in business-to-business marketing and CRM services like Salesforce and Mailgun. Like the overall email volume, the email volume by ASN also follows the Pareto principle, where a few large providers are responsible for the sending of the large majority of emails. Specifically, the top twenty percent, 8 ASNs, produced 89.35% of the total email volume. Figure 3 shows the evidence that most of the email marketing flow originates from a few large, third-party providers.
@@
However, understanding which companies use which email marketing service, or whether they use a third-party service at all, is not straightforward. Some companies send emails originating from their own registered ASN such as wish.com (Wish is the “doing-business-as” name for ContextLogic), Alibaba Group, LinkedIn, and Adobe (Adobe acquired Omniture as part of their web analytics platform). Others employ specialized third-parties, chiefly Salesforce, Mailgun, Sendgrid, and Sparkpost. A third group had emails originate from cloud service providers like AWS, GCP, and Rackspace. For those that originate from cloud service providers, using only the IP addresses we cannot ascertain what entity is managing the email campaigns at the application layer.
Sender Reputation. While the domain-level analysis confirms authenticity and shows no evidence of unauthorized or unknown senders, the infrastructure-level analysis reveals a potentially more ambiguous trust environment. To deepen our understanding of the quality and trustworthiness of these sending infrastructures, we conducted an IP reputation analysis. We cross-referenced each unique email origination IP against a third-party IP reputation database (AbuseIPDB, [n. d.]), which aggregates historical spam reports and other abuse metrics. In total, there were 612 spam reports associated with 84 unique IP addresses belonging to 11 ASNs, corresponding to 29 unique companies that were in our study. This allows us to assess the prevalence of reported spam activity associated with the IP ranges, from which we received emails, used by various ASNs.
By coupling the ASN-level volume analysis with the historical IP reputation data, we found that even well-known brands may rely, potentially unknowingly, on email infrastructure that carries varying levels of reported spam activity. The treemap visualization in Figure 4 arranges ASNs hierarchically, with each company’s sending domain nested beneath the ASN that serviced it. The area of each box corresponds to the total number of spam reports for the associated IPs, while the grayscale intensity indicates the magnitude of these reports.
This hierarchical perspective reveals patterns and raises important questions for both research and policy. For instance, some large ASNs have a relatively high volume of reported spam across multiple well-known clients. In such cases, the individual brand may not be involved in illicit or spammy behaviors, yet they inherit the reputation challenges posed by their chosen infrastructure providers. Or, the community-writ-large will deem certain messages as spam and report them as abuse. These findings highlight that brand trust does not necessarily translate into a pristine underlying email infrastructure. Rather, the operational practices, vetting procedures, and network hygiene of intermediary email service providers matter a great deal.
Ultimately, these findings show that the email ecosystem is more complex and federated than it may appear on the surface. While the messages themselves may appear genuine and from the expected sender, the underlying infrastructure often involves multiple intermediaries and networks of varying reputations.
For end-users, this complexity raises concerns about data stewardship and privacy. When personal information and communication preferences flow through third-party service providers, some of which may have been associated with spam-related activity, users face heightened uncertainty about how their data is handled, who can access it, and to what end. This underscores the need for greater transparency and more privacy-centric policies from both the companies authorizing these messages and the infrastructure providers enabling their delivery.
Figure 4. Hierarchical treemap of community reported spam on sender IPs, organized by ASN and associated company domains. Lighter shades indicate higher counts of spam reports, and vice versa.
IP Hopping and Spam Reports. To investigate whether companies associated with higher spam reports are more likely to use a larger number of IPs, a behavior that could suggest “IP hopping” to evade reputation-based spam filters, we examined the correlation between the number of unique IPs per company and the total spam reports associated with those IPs.
[…]
This result highlights that companies linked to higher spam reports may be employing a strategy of cycling through multiple IP addresses to distribute the reputational impact across their infrastructure. Such IP hopping practices can dilute the effects of blacklistings or abuse reports on any single IP address (Ramachandran et al., 2007), allowing these companies to maintain email delivery volumes despite the reputational damage incurred by individual IPs.